a cell divides every 20 minutes how many cells would be produced from one original cell at the end of two hours
1
Expert's answer
2016-09-22T11:27:04-0400
GP is An = a(r)^n
r - the number of progenies that are 2 (since one cell gives 2 cells)
a - the start number of cells, here is 1.
n - the number of divisions a cell can undergo in a given time period, which equals, (2 hours · 60 minutes)/ 20 minutes) = 6 times.
An - the number of cells after 2 hours.
An = 1 x 2^6 = 64 cells.